# Open-world Machine Learning: A Review and New Outlooks

### v2

@@ 2.3 Overall Challenges of Open-world Learning
As illustrated in Figure 1 and Figure 2, open-world learning involves performing unknown rejection, novel class discovery, and class-incremental learning sequentially and periodically. The core challenge is enabling the above process to automatically proceed through the interactions between the model and the open environments without relying on human engineers [26]. Unfortunately, under the close-world assumption [7], a model is overconfident and can hardly be aware of the unknown. Specifically, from the perspective of representation learning, the model is trained only on the current dataset with data-driven optimization, the learned representations are task-specific and less generic; from the perspective of classifier learning, current discriminative classifiers leave little room for the unknown, making it hard to characterize, discover and adapt to novelty. Consequently, examples from unknown classes would be mapped to the region of known classes, leading to catastrophic forgetting of previous knowledge in the latter incremental learning process.
3 Unknown Rejection
@@
Unknown rejection is the first step towards open-world machine learning, which is also a fundamental ability of the classifier in the open-world. Considering that a training set comprises of $K$ classes, i.e., $\mathcal{D}_{\text{train}}=\mathcal{D}_{\text{in}}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}\subset\mathcal{X}\times\mathcal{Y}_{\text{in}}$, where $\mathcal{Y}_{\text{in}}=\{1,2,\cdots,K\}$. Once trained on $\mathcal{D}_{\text{in}}$, models are supposed to reject unknown samples from classes outside of $\mathcal{Y}_{\text{in}}$, rather than classifying it irresponsibly into one of $K$ categories. The philosophy behind unknown rejection is that when encountering unfamiliar knowledge, it is essential to bravely acknowledge the limitations rather than feign understanding and give arbitrary answers.
@@
Many efforts have been made to enhance the unknown rejection ability of machine learning systems. There are multiple research areas related to unknown rejection, such as anomaly detection [27, 28], out-of-distribution (OOD) detection [29, 30], and open-set recognition [31, 32] (OSR). Among them, anomaly detection is widely used in early work, while OOD detection and OSR are more often used in recent works. The differences between OOD detection and OSR lie in two aspects. First, in OOD detection, the semantic distance between the OOD data and in-distribution (ID) data is relatively larger, e.g., the OOD dataset is usually completely unrelated. While in OSR, a subset of classes from a dataset is viewed as ID and other classes from the same dataset are viewed as OOD data. Second, OOD detection mainly concentrates on distinguishing OOD samples from ID samples, whereas OSR also evaluates closed-world classification performance on known classes. We overview the recent advances in OOD detection and OSR in the remaining part of this section, and different types of methods are illustrated in Figure 3.
@@
Key Challenges. The inherent challenge of unknown rejection is that OOD samples are agnostic during training. Even though some methods [33, 34] resort to auxiliary outlier OOD samples, it is impossible to exhaustively encompass all OOD distributions. As a consequence, models need to reverse some open space for potential unknown samples to reduce the open space risk [31, 32]. Secondly, when the semantic similarity [35] between OOD and ID samples is minimal, their separability decreases, resulting in a decline in rejection capability. Such samples are referred to as near-ood samples in the literature [36]. Furthermore, some issues could impact the rejection of unknown samples, including covariate shifts [37] and spurious correlations [38], for example, ID samples from another distribution domain might be mistakenly detected as OOD samples, and spurious correlations learned from training data could mislead the models to focus on irrelevant features and areas.
Figure 3: Illustrations of different types of methods of (A) OOD detection and (B) OSR. In OOD detection, methods are divided according to training and inference strategies while modeling perspectives in OSR. In OSR, “loss D”, “loss G” and “loss H” denote loss functions for discriminative, generative and hybrid models, respectively.
3.1 Out-of-distribution Detection
@@
OOD detection focuses on whether a test example is outside the distribution of the training data and aims to improve the separability between ID and OOD samples. The most widely used metric for OOD detection is the Area Under the Receiver Operating Characteristic curve [39] (AUROC), which is a threshold-independent metric that can be interpreted as the probability that a positive example (ID) is assigned a higher prediction score than a negative example. In addition, there are many other metrics like TNR-$95\%$-TPR, AUPR-In and AUPR-Out [10] to imply how the in-distribution and OOD samples are separated. Technically, the common practice is to assign each sample with a score function $S(\mathbf{x})$ indicating the otherness of each sample. If the score is above a pre-defined threshold $\delta$, then the sample is recognized as ID, otherwise detected as OOD and rejected. OOD detection methods can be categorized into three classes [36, 40] based on different training and inference techniques.
@@
Post-hoc Inference Methods. One natural and simple way is to design appropriate score functions $S(\mathbf{x})$ that have the most significant separability between ID and OOD samples. Post-hoc methods are simply implemented on the trained models, without interfering with the training phase, as a result, they are orthogonal to the training-stage methods discussed below. Many characteristics of the trained models could be designed as score functions, such as posterior probabilities of the classification head, activations and dynamics in the feature space. Early works design $S(\mathbf{x})$ based on probabilities, i.e., $p(c|\mathbf{x})=\frac{\exp(\mathbf{z}_{c}(x))}{\sum_{j=1}^{K}\exp(\mathbf{z}_{j}(x))}$ where $\mathbf{z}_{c}(x)$ denotes the logits of sample $\mathbf{x}$. Hendrycks et al. [10] proposed a simple score function, namely maximum softmax probability (MSP), i.e. $S(\mathbf{x})=\max_{c}p(c|\mathbf{x})$, which is regarded a baseline method in the literature. MSP could be enhanced by input perturbation which aims to increase the softmax score for the true label and temperature scaling [41]. Apart from softmax probabilities, energy score [42] derived from energy-based models [43] is less susceptible to the overconfidence issue of classifiers, achieving better OOD detection abilities. Similar to probabilities, logits also serve as the belongingness of each class without the normalization process in MSP, and maximum logit score (MLS) [44] is reported better than MSP score. Instead of the probability distribution, characteristics in the feature space could also be exploited for OOD detection. Lee et al. [45] designed the Mahalanobis distance-based confidence score by modeling layer-wise Gaussian distribution of features, which could be substituted by flow-based [46] generative modeling for more flexible expressivity [47]. These methods still impose certain assumptions on the distribution of the feature spaces, which limits the performance to some extent. By contrast, non-parametric nearest-neighbor distance [48] is more general and flexible without relying on any distributional assumptions. Activations in the penultimate layer [49] could also be considered. The methods above design $S(\mathbf{x})$ from only one of the two aspects. We could also combine the above two perspectives [50], namely posterior probabilities and feature embeddings, for a more remarkable separation between ID and OOD samples.
@@
Training-stage Methods. Training-based methods directly explore the training of models with inherent capabilities to discriminate between ID and OOD samples. This line of works are complementary to the post-hoc methods dedicated to designing OOD scores, they could be combined for better OOD detection performance. In general, advanced training tricks including self-supervised learning [51] and normalization [52], provide outstanding OOD detectors. Self-supervised learning, like rotation prediction [53] and contrastive learning [54], could help improve robustness and uncertainty estimation, which also greatly boost OOD detection [51, 55] with alleviated overconfidence issues. Another technique namely logit normalization [52] helps to mitigate the overconfidence issue for both ID and OOD samples. Models could also be encouraged for greater inter-class separation and intra-class compactness [56] in the feature space, which leaves more room for open space and helps separate ID and OOD samples. Recently, flat minima has been demonstrated to be a good optimization objective for the detection of misclassified examples from known classes and OOD examples from unknown classes [15]. When scaling OOD detection to large-scale datasets, a key insight is to decompose the large semantic space into many small groups and train models with group-based learning [57].
Outlier-aided Methods. It is also intuitive for better OOD performance by directly training models on auxiliary OOD samples $\mathcal{D}_{\text{out}}$ and maximizing the uncertainty on them. Outlier exposure [33] (OE) is a pioneer work to explicitly require and learn on the auxiliary outliers. Specifically, OE proposes to maximize the uncertainty on the outliers. The loss function of OE is $\mathcal{L}=\mathbb{E}_{(\mathbf{x},y)\in\mathcal{D}_{\text{in}}}\mathcal{L}_{ce}(y,f(\mathbf{x})+\lambda\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\text{out}}}\mathcal{L}_{ce}(\mathcal{U},f(\mathbf{x}))$, where $f(\mathbf{x})$ is the predictive probability, $\mathcal{L}_{ce}$ is the cross-entropy loss, and $\mathcal{U}$ is the uniform distribution. Practically, OE performs to minimize the cross-entropy between the uniform distribution and the model’s predictions to encourage uncertain predictions. The authors [33] found that once learned on $\mathcal{D}_{\text{out}}$, the model could generalize to detect unseen OOD samples and anomalies and achieves remarkable OOD detection performance. Recently, Zhu et al. [58] revealed that popular OOD detection methods like OE often make it harder to detect misclassified examples from known classes, and proposed a unified misclassification and OOD detection approach named OpenMix.
3.2 Open-Set Recognition
@@
OSR [31, 32] aims to recognize known classes (ID) while detecting unknown classes (OOD) at the same time. In the literature of OSR, ID and OOD are conventionally drawn from different class sets within the same dataset. Early OSR works are mainly based on support vector machine [31], nearest class mean classifiers [59] and extreme value machine [60]. Those traditional machine learning-based OSR approaches have been well reviewed and discussed [32]. In recent years, various deep learning-based OSR methods have been developed, and we mainly focus on them in this paper. The methods could be divided into three types from the modeling perspective: discriminative, generative and hybrid models.
@@
Discriminative Models. Discriminative models perform relatively well in classification tasks, it is natural to employ them to handle OSR, i.e., classify among known classes and reject unknown classes. One direct and intuitive method is to leave room for the modeling of unknown classes. A pioneer work OpenMax [61] re-calibrates the logits of top classes from known to assign the probability to the appended class zero representing unknown classes. OpenMax [61] resorts to extreme value theorem and utilizes Weibull distribution to compute the re-calibration weights. Dummy classifiers [62] for unknown samples could mimic the novel classes as data placeholders and augment the closed-set classifier with classifier placeholders. From another perspective, similar to OOD detection, models could also be trained on both known and auxiliary unknown classes to behave differently on knowns and unknowns and improve the ability to distinguish between them. Dhamija et al. [63] leveraged samples from background classes as unknown samples and proposed two novel loss functions to make ID and OOD well-separated in both probability space and feature space. DTL [64] proposes membership loss to substitute cross-entropy loss for class-wise probability modeling, and also relies on outlier data to train a parallel network for providing global negative filtering. The methods discussed above mainly employ the output of classifiers as a proxy to discern ID and OOD samples. Besides, auto-encoders are proven good indicators for OSR, the fundamental principle is that models are trained on known samples to minimize reconstruction errors and will produce relatively higher reconstruction errors on unknown samples. C2AE [65] divides OSR into closed-set classification and open-set identification, and the test sample is rejected as unknown if the reconstruction error is below the threshold.
Generative Models. Generative models provide different perspectives from discriminative models and could be utilized mainly in two ways, i.e., one is generative and probabilistic modeling of known and unknown classes, and the other is to explicitly generate fake unknown samples for greater separation between known and unknown classes. From the perspective of generative modeling, the basic idea is to model posterior distributions in the latent space to approximate conditional Gaussian distributions [66] and the probability of test samples could be used for unknown rejection. However, directly modeling the distribution is difficult, and most existing methods employ off-the-shelf generative models to generate pseudo-unknown samples to augment classifiers. For example, G-OpenMax [67] generates unknown samples with GANs [68] and fine-tunes the model with OpenMax [61] loss. OpenGAN [69] adopts an adversarial training framework to generate pseudo-unknown samples on the deep feature space. By further taking the difficulty [70] of the generated unknown samples into consideration, models could grasp fine-grained concepts of unknowns for results. Besides, without using explicit generative models, pseudo-unknown samples could be generated by optimizing a latent vector to have a small distance to training samples but produce low confidence on known categories [71]. These samples could be viewed as difficult open-space samples which are similar to known samples. A $(K+1)$-way classifier is trained on the combination of training samples and the generated counterfactual samples.
Hybrid Models. Discriminative and generative models are complementary to each other, and both have their specific advantages. In principle, discriminative models perform better on classification tasks than generative models. On the other hand, generative models could formalize the distribution of known and unknown samples, and additionally generate fake samples to augment discriminative models. As a consequence, combining both discriminative and generative models into hybrid models results in enormous potential for performing OSR tasks. Zhang et al. [72] proposed to train a flow-based generative model [73] and an inlier classifier on a shared feature space of the encoder. The flow-based model and the classifier are targeted to detect outliers and classify inliers respectively without intervention. The flow-based model is trained on the feature space using the maximum likelihood loss function, which avoids the issue of assigning OOD with high probability [74] when learned in the input pixel space. Yang et al. [75, 76] creatively proposed convolutional prototype networks (CPN) for OSR with learnable class-wise prototypes serving as a classifier, which is different from conventional linear classifiers. The prototypical classifier could be viewed as Gaussian generative modeling for each class, and leaves more room for open set samples in the feature space. CPN is trained on both discriminative and generative loss functions to encourage inter-class separation and intra-class compactness respectively, resulting in less open set risk and remarkable OSR performance. ARPL [77, 78] takes the similar spirit of CPN [76], but models class-wise reciprocal points representing the extra-class space, i.e., modeling the open-space of each category. ARPL [78] additionally employs adversarial enhancement to generate diverse confusing samples to further reduce the open space risk caused by confusing samples. The class-wise points of CPN [76] and ARPL [78] could be generalized to class-wise continuous manifolds [79], which are constructed by class-wise auto-encoders upon the feature space. Each class-wise manifold could be viewed as an infinite number of prototype/reciprocal points, and thus more flexible and generalizable.
3.3 Analysis of Unknown Rejection

### v3

@@ II-C Overall Challenges of Open-world Learning
As illustrated in Fig. 2 and Fig. 3, open-world learning involves performing unknown rejection, novel class discovery, and continual learning sequentially and periodically. The core challenge is enabling the above process to automatically proceed through the interactions between the model and the open environments without relying on human engineers [250]. Unfortunately, under the closed-world assumption [9], a model is overconfident and can hardly be aware of the unknown. Specifically, from the perspective of representation learning, the model is trained only on the current dataset with data-driven optimization, the learned representations are task-specific and less generic; from the perspective of classifier learning, current discriminative classifiers leave little room for the unknown, making it hard to characterize, discover and adapt to novelty. Consequently, examples from unknown classes would be mapped to the region of known classes, leading to catastrophic forgetting of previous knowledge in the latter continual learning process.
III Unknown Rejection
III-A Problem Formulation and Key Challenges
@@
1) Problem Formulation. Unknown rejection is the first step towards open-world machine learning, which is also a fundamental ability of the classifier in the open-world. Considering that a training set comprises of $K$ classes, i.e., $\mathcal{D}_{\text{train}}=\mathcal{D}_{\text{in}}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}\subset\mathcal{X}\times\mathcal{Y}_{\text{in}}$, where $\mathcal{Y}_{\text{in}}=\{1,2,\cdots,K\}$. Once trained on $\mathcal{D}_{\text{in}}$, models are supposed to reject unknown samples from classes outside of $\mathcal{Y}_{\text{in}}$, rather than classifying them irresponsibly into one of $K$ categories. The philosophy behind unknown rejection is that when encountering unfamiliar knowledge, it is essential to bravely acknowledge the limitations rather than feign understanding and give arbitrary answers.
@@
Many efforts have been made to enhance the unknown rejection ability of machine learning systems. There are multiple research areas related to unknown rejection, such as anomaly detection [251, 252], out-of-distribution (OOD) detection [253, 254], and open-set recognition [255, 256] (OSR). Among them, anomaly detection is widely used in early work, while OOD detection and OSR are more often used in recent studies. The differences between OOD detection and OSR lie in two aspects. First, in OOD detection, the semantic distance between the OOD data and in-distribution (ID) data is relatively larger, e.g., the OOD dataset is usually completely unrelated. While in OSR, a subset of classes from a dataset is viewed as ID, and other classes from the same dataset are viewed as OOD data. Second, OOD detection mainly concentrates on distinguishing OOD samples from ID samples, whereas OSR also evaluates closed-world classification performance on known classes. We overview the recent advances in OOD detection and OSR in the remaining part of this section, and different types of methods are summarized in Fig. 4 and illustrated in Fig. 5.
@@
2) Key Challenges. The inherent challenge of unknown rejection is that OOD samples are agnostic during training. Even though some methods [32, 257] resort to auxiliary outlier OOD samples, it is impossible to encompass all OOD distributions exhaustively. As a consequence, models need to reserve some open space for potential unknown samples to reduce the open space risk [255, 256]. Secondly, when the semantic similarity [258] between OOD and ID samples is minimal, their separability decreases, resulting in a decline in rejection capability. Such samples are referred to as near-ood samples in the literature [259]. Furthermore, some issues could impact the rejection of unknown samples, including covariate shifts [260] and spurious correlations [261], for example, ID samples from another distribution domain might be mistakenly detected as OOD samples, and spurious correlations learned from training data could mislead the models to focus on irrelevant features and areas. In this section, we will introduce two types of tasks related to rejection, i.e., OOD detection (Sec. III-B) and Open-set Recognition (Sec. III-C), respectively, as well as their representative methods and comparative results. Then, the theoretical advance of unknown rejection is summarized (Sec. III-D). Finally, evaluation results of unknown rejection are presented (Sec. III-E).
Figure 4: The evolutionary tree of unknown rejection methods.
III-B Out-of-distribution Detection
@@
OOD detection [253, 254, 35] focuses on whether a test example is outside the distribution of the training data and aims to improve the separability between ID and OOD samples. Technically, the common practice is to assign each sample a score function $S(\bm{x})$, indicating the otherness of each sample. If the score is above a pre-defined threshold $\delta$, then the sample is recognized as ID, otherwise it is detected as OOD and rejected. This formula could be written in Eq. (1). OOD detection methods can be categorized into three classes [259, 262] based on different training and inference techniques.
$g(\bm{x})=\left\{\begin{aligned} &\text{in-distribution},\quad&\text{if}\ S(\bm{x})\geq\delta,\\ &\text{out-of-distribution},\quad&\text{if}\ S(\bm{x})<\delta.\end{aligned}\right.$ (1)
@@
1) Post-hoc Inference Methods. One natural and simple way is to design appropriate score function $S(\bm{x})$ that has the most significant separability between ID and OOD samples. Post-hoc methods are simply implemented on the trained models, without interfering with the training phase, as a result, they are orthogonal to the training-stage methods discussed below. Many characteristics of the trained models could be designed as score functions, such as posterior probabilities of the classification head, activations and dynamics in the feature space. Early works design $S(\bm{x})$ based on probabilities, i.e., $p(c|\bm{x})=\frac{\exp(\bm{z}_{c}(x))}{\sum_{j=1}^{K}\exp(\bm{z}_{j}(x))}$ where $\bm{z}_{c}(x)$ denotes the logits of sample $\bm{x}$. Hendrycks et al. [10] proposed a simple score function, namely maximum softmax probability (MSP), i.e., $S(\bm{x})=\max_{c}p(c|\bm{x})$, which is regarded as a baseline method in the literature. MSP could be enhanced by input perturbation, which aims to increase the softmax score for the true label and temperature scaling [11]. Apart from softmax probabilities, energy score [12] derived from energy-based models [263] is less susceptible to the overconfidence issue of classifiers, achieving better OOD detection abilities. Similar to probabilities, logits also serve as the belongingness of each class without the normalization process in MSP, and the maximum logit score (Maxlogits) [22] is reported better than the MSP score. Instead of the probability distribution, characteristics in the feature space could also be exploited for OOD detection. Lee et al. [13] designed the Mahalanobis distance-based confidence score by modeling layer-wise Gaussian distribution of features, which could be substituted by flow-based [264] generative modeling for more flexible expressivity [265]. These methods still impose certain assumptions on the distribution of the feature spaces, which limits the performance to some extent. By contrast, non-parametric nearest-neighbor distance (KNN) [16] is more general and flexible without relying on any distributional assumptions. Activations in the penultimate layer [266] could also be considered. The methods above design $S(\mathbf{x})$ from only one of the two aspects. We could also combine the above two perspectives [14], namely posterior probabilities and feature embeddings, for a more remarkable separation between ID and OOD samples. Virtual-logit matching (ViM) [14] combines the class-agnostic information in the feature space and the class-dependent information in logits. Recently, activation shaping has been demonstrated to be quite effective for OOD detection. For example, ASH [19] enhances OOD detection performance remarkably by pruning and lightly adjusting a large portion of late-layer activations during inference. Similarly, by scaling network activations, SCALE [20] achieves state-of-the-art performance without compromising in-distribution accuracy.
Figure 5: Illustrations of different types of methods of OOD detection (first row) and Open-Set Recognition (OSR, second row). In OOD detection, methods are divided into three types, i.e., (a), (b) and (c), according to training and inference strategies. For OSR, methods are also divided into three types, i.e., (d), (e) and (f), regarding modeling perspectives in OSR. In OSR, “loss D”, “loss G” and “loss H” denote loss functions for discriminative, generative and hybrid models, respectively. Here, $f$ and $g$ denote the feature extractor and generator, respectively.
@@
2) Training-stage Methods. Training-based methods directly explore the training of models with inherent capabilities to discriminate between ID and OOD samples. This line of works are complementary to the post-hoc methods dedicated to designing OOD scores, they could be combined for better OOD detection performance. In general, advanced training tricks, including self-supervised learning [23] and normalization [24], provide outstanding OOD detectors. Self-supervised learning, like rotation prediction [267] and contrastive learning [268], could help improve robustness and uncertainty estimation, which also greatly boosts OOD detection [23, 269] with alleviated overconfidence issues. Another technique, namely logit normalization [24], helps to mitigate the overconfidence issue for both ID and OOD samples. Models could also be encouraged for greater inter-class separation and intra-class compactness [270] in the feature space, which leaves more room for open space and helps separate ID and OOD samples. Recently, flat minima has been demonstrated to be a good optimization objective for the detection of misclassified examples from known classes and OOD examples from unknown classes [25]. Recently, Chen et al. [30] revealed that training with large-scale synthetic images from generative models can effectively eliminate overfitting in OOD detection, enabling small models like ResNet-18 [271] trained on generated data to rival ViT [272] pretrained on massive real datasets. When scaling OOD detection to large-scale datasets, a key insight is to decompose the large semantic space into many small groups and train models with group-based learning [21]. TrustLoRA [273] introduces the low-rank adaptation technique into unknown rejection and unifies rejection under both covariate and semantic shifts via reliability consolidation.
3) Outlier-aided Methods. It is also intuitive to directly train models on auxiliary OOD samples $\mathcal{D}_{\text{out}}$. By maximizing their uncertainty, the model tends to assign lower confidence to OOD samples, which helps increase the separability between OOD and ID samples. Outlier exposure [32] (OE) is a pioneering work to leverage auxiliary outliers during training. Specifically, OE proposes to maximize the uncertainty on the outliers. The loss function of OE is:
$\mathcal{L}=\mathbb{E}_{(\bm{x},y)\in\mathcal{D}_{\text{in}}}\mathcal{L}_{ce}(y,f(\bm{x})+\lambda\mathbb{E}_{\bm{x}\in\mathcal{D}_{\text{out}}}\mathcal{L}_{ce}(\mathcal{U},f(\bm{x})),$ (2)
where $f(\bm{x})$ is the predictive probability, $\mathcal{L}_{ce}(\cdot,\cdot)$ is the cross-entropy loss, and $\mathcal{U}$ is the uniform distribution. Practically, OE aims to minimize the cross-entropy between the uniform distribution and the model’s predictions to encourage uncertain predictions on these outlier samples. The authors [32] found that once learned on $\mathcal{D}_{\text{out}}$, the model could generalize to detect unseen OOD samples and anomalies and achieve remarkable OOD detection performance. Zhu et al. [33] revealed that popular OOD detection methods like OE often make it harder to detect misclassified examples from known classes, and proposed a unified misclassification and OOD detection approach named OpenMix. Besides, taking the spirit of MixUp [274], mixing ID training samples with auxiliary outliers [257] could expand the coverage of different OOD granularities and enhance the performance of OE. Virtual Outlier Synthesis (VOS) [34] models the penultimate feature embeddings of each class as a multivariate Gaussian, then samples virtual outliers from the low‐likelihood regions of these class‐conditional distributions and uses them to tighten the ID/OOD decision boundary. Ma et al. [35] extend the setting of OOD detection to the data-efficient scenarios [275, 276, 277], and propose to synthesize pseudo-outliers using ID data, which could also improve the detection performance without the requirement of real outliers. Recent studies [36, 37, 38, 39] focus on few-shot OOD detection of vision–language models with pseudo-outliers mining. For example, LoCoOp [36] leverages local image features as negative OOD examples during prompt training to regularize and purify the context embeddings, thereby removing ID-irrelevant nuisances and enhancing ID/OOD separation. LocalPrompt [39] uses prompt guided negative augmentation to inject local outlier cues to improve the performance of rejecting unknown rejection.
III-C Open-Set Recognition
@@
OSR [255, 256] aims to recognize known in-distribution classes while detecting unknown OOD classes at the same time. In the literature of OSR, ID and OOD are conventionally drawn from different class sets within the same dataset. Early OSR works are mainly based on support vector machine [255], nearest class mean classifiers [278], and extreme value machine [279]. Those traditional OSR approaches have been well reviewed and discussed [256]. In recent years, various deep learning-based OSR methods have been developed, and we mainly focus on them in this paper. The methods could be divided into three types from the modeling perspective: discriminative, generative and hybrid models.
@@
1) Discriminative Models. Discriminative models perform relatively well in classification tasks, it is natural to employ them to handle OSR, i.e., classify among known classes and reject unknown classes. One direct and intuitive method is to leave room for the modeling of unknown classes. A pioneer work, OpenMax [41] recalibrates the logits of top classes from known to assign the probability to the appended class zero representing unknown classes. OpenMax [41] resorts to the extreme value theorem and utilizes the Weibull distribution to compute the re-calibration weights. Dummy classifiers [42] for unknown samples could mimic the novel classes as data placeholders and augment the closed-set classifier with classifier placeholders. From another perspective, similar to OOD detection, models could also be trained on both known and auxiliary unknown classes to behave differently on knowns and unknowns and improve the ability to distinguish between them. Dhamija et al. [280] leveraged samples from background classes as unknown samples and proposed two novel loss functions to make ID and OOD well-separated in both probability space and feature space. DTL [281] proposes membership loss to substitute cross-entropy loss for class-wise probability modeling, and also relies on outlier data to train a parallel network for providing global negative filtering. The methods discussed above mainly employ the output of classifiers as a proxy to discern ID and OOD samples. Besides, autoencoders [282, 283] are proven to be good indicators for OSR. The fundamental principle is that models are trained on known samples to minimize reconstruction errors and will produce relatively higher reconstruction errors on unknown samples. C2AE [43] divides OSR into closed-set classification and open-set identification, and the test sample is rejected as unknown if the reconstruction error is below the threshold. The encoder is first trained on known training samples to learn the classifier. The second stage is conditional decoder training to reconstruct inputs conditioned on class identity. C2AE [43] employs EVT to obtain the threshold for unknown identification. The test sample is rejected as unknown if the reconstruction error is below the threshold. Furthermore, maintaining hierarchical reconstructions [284] with various levels of abstraction brings about more stable results.